OpenStudy (anonymous):

for f(x)=3x^2+7x+6, find all values of a for which f(a)=4. a=?? (use commas to seperate answers as needed)

OpenStudy (anonymous):

Ok, sub a in for x and set the function equal to 4. Solve for a from there.

OpenStudy (anonymous):

ok u lost me lol

OpenStudy (anonymous):

Ok, you're starting with f(x), but then it asks about f(a), that means x=a, so change that first, because that's what it's asking for.

OpenStudy (anonymous):

so it will be f(a)=3a^2+7(a)+6=4 ?

OpenStudy (anonymous):

f(a)=4, so after you do that substitution, set the whole function of a = 4, then start solving.

OpenStudy (anonymous):

Yes, exactly. See what to do next?

OpenStudy (anonymous):

3(4)^2+7(4)+6=4 right?

OpenStudy (anonymous):

No, you don't know if a is 4 yet. You know f(a)=4, which means whatever you find for a, when you put that value in it'll make the whole thing equal to 4.

OpenStudy (anonymous):

ok so what is my next step

OpenStudy (anonymous):

You got your equation, 3a^2+7(a)+6=4 Quadratic equations like this one are easiest to solve when set equal to zero, so if you subtract 4 from each side, the right side goes to zero.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

then?

OpenStudy (anonymous):

It's factorable . . .

OpenStudy (anonymous):

so i factor 3a^2+7a+6=4 ?

OpenStudy (anonymous):

But then the quadratic formula always works, so you could use that too.

OpenStudy (anonymous):

Subtract 4 from each side and then factor. Get it equal to zero first.

OpenStudy (anonymous):

ok i got it..

OpenStudy (anonymous):

I knew you would. :-)